### v2

# Learning Nonlinear Finite Element Solution Operators using Multilayer Perceptrons and Energy Minimization

Abstract
We develop and evaluate a method for learning solution operators to nonlinear problems governed by partial differential equations (PDEs). The approach is based on a finite element discretization and aims at representing the solution operator by a multilayer perceptron (MLP) that takes problem data variables as input and gives a prediction of the finite element solution as output. The variables will typically correspond to parameters in a parametrization of input data such as boundary conditions, coefficients, and right-hand sides. The output will be an approximation of the corresponding finite element solution, thus enabling support and enhancement by the standard finite element method (FEM) both theoretically and practically. The loss function is most often an energy functional and we formulate efficient parallelizable training algorithms based on assembling the energy locally on each element. For large problems, the learning process can be made more efficient by using only a small fraction of randomly chosen elements in the mesh in each iteration. The approach is evaluated on several relevant test cases, where learning the finite element solution operator turns out to be beneficial, both in its own right but also by combination with standard FEM theory and software.
1 Introduction
In recent years, there has been a lot of interest in using machine learning methods in applied mathematics and scientific computing, see for example [31]. In this work, we focus on methods that approximate the solution operator to parameterized partial differential equations (PDEs). We emphasize that the aim is not to learn the solution to a single PDE problem, but rather to a family of the same type of PDE problems, by learning the mapping from a low-dimensional parameter space into the solution space. As the solution space is typically infinite-dimensional we consider discrete spaces associated with standard numerical method for PDEs. We thus aim at learning well-known approximations instead of the exact solution. The idea is that the standard numerical method could support and enhance the machine learning one. Here we consider the finite element method (FEM) and present some examples of beneficial combination with the machine learning framework.
There are several approaches to learning solution operators or solutions to PDEs in the literature, see for example [19] where neural operators on a general integral form together with some concrete examples are presented. The different approaches can be divided into two main branches; data-driven or physics-informed. The data-driven methods are based on (typically pre-computed) data, such as, e.g., the neural operators in [21], Fourier neural operators [22], DeepONets [24], and the random feature model [26]. The physics-informed methods typically penalize the network to satisfy a PDE or other physical laws by designing an appropriate loss function, e.g., an energy-based loss function such as in the Deep-Ritz method [11] or a loss function based on the strong residual as for Physics-informed neural networks (PINNs) [30]. These methods do not require any data in general. Typically, the focus of physics-informed approaches has not been on learning parameterized problems, but mainly the solution to a single PDE problem, often with collocation type methods such as [11, 30, 7, 25, 36]. Recently however, works on learning parameterized PDEs seem to have gained more traction since they allow for a single network to predict the solutions to a class of related problems, see, e.g., [35, 13, 14, 17], or [10, 9] where PINNs is extended to parameterized settings, and [38] in which networks are trained to parameterize the Green function. Several parameterized-PDE approaches also take into account the variational formulation in some manner: [3] in a loss function related to the error in the norm induced by the variational formulation, [37] by basing the the loss function on the weak residual, and [29] instead in the network architecture. A work similar to the current is [17]. There, the map from a parameter space into a finite element space is also learned and some analysis is presented together with examples. However, there a field-to-field map is considered with convolutional neural networks (a natural choice for such a map) and both the analysis and examples are for linear problems. Here, we consider a tuple-to-field map with MLPs and both analysis and examples include nonlinearities. Another similar work is [13], where so called mesh-informed neural networks (MINNs) are used to map parameter tuples into a finite element space. There the loss function is based on the strong residual, not energy as here, also there is no analysis but several impressive examples are presented.
In this paper we consider a data-free physics-informed method for operator learning of parameterized PDEs. The method is actually a special case of the more general approach presented in [34] and it has already been applied to inverse problems in [8]. However, in the former, the method is barely noticeable in the more general framework presented there. In the latter, the focus lies on the application rather than on the method. Here, we therefore put full attention on the method itself, develop and present it in a general form along with some analysis and examples.
We construct a neural network that learns a family of solutions by utilizing a loss function based on energy minimization that penalizes the network to satisfy the PDE. This means that the computationally costly step of generating training data is avoided. The energy minimization approach has earlier been considered in neural network contexts, e.g., [11, 33, 20, 12] and [34] for finding low-dimensional models of physical motion and [32] for applications in soft tissue modeling. In this work we consider a method that can be applied to any (nonlinear) problem with a given energy functional. The problem is first discretized using a finite element method and a multilayer perceptron (MLP) neural network is trained to learn the corresponding approximate solution operator. The input to the neural network is typically parameters to the problem which define the data on, e.g., the boundary, the coefficients, or the right-hand side. The output of the network is the finite element degrees of freedom (DOFs), which for, e.g., classical P1 elements are the function values at the nodes of the finite element mesh.
The general idea behind learning finite element solution operators is that regardless of machine learning method, a predicted PDE solution will be on a discrete form. It might therefore be reasonable to consider a discrete form related to a well-studied established method, rather than some seemingly arbitrary form specific to the learning approach. A motivation being that the established method might piggyback the machine learning one to enhance the final product. The established method we consider here is FEM due to its ubiquity, flexibilty, and substantial theory, but the general idea applies to any standard numerical method together with any machine learning framework. See Figure 1 for an illustration representing this general idea. The operator networks we consider thus map into a finite element space $V_{h}$ which allows this machine learning approach for PDEs to smoothly be combined with the powerful finite element method. This has a couple of advantages. Firstly, in analyzing the approximation error of the learned solution, tools from finite element theory may be incorporated to provide better error estimates. Secondly, existing finite element software may be used to improve the learned results. In this work we provide examples of both of these advantages.
Furthermore, we provide a naturally parallelizable algorithm for computing the energy in each iteration, which is necessary for efficient training on GPUs. We also suggest to use random batches of elements in the mesh, which has potential to be useful in large problems with many degrees of freedom.
The paper is organized as follows. In Section 2 we present the general setting by introducing abstract spaces and energy functionals. In Section 3 we describe the neural network approach and define appropriate loss functions based on the energy. Finally, in Section 4 we present several numerical examples where operator learning turns out to be useful.
$u$
$u_{\theta}$
$\mathbb{R}^{n_{\text{output}}}$
$u$
$u_{h,\theta}$
$V_{h}$
$u_{h}$
Figure 1: Left: Simplified default approach to learning PDE-solution $u$. Right: Approach of instead learning the corresponding finite element solution $u_{h}$.
2 General setting and energy functionals
Consider a general nonlinear PDE which admits an energy functional $E:V\to\mathbb{R}$ for a suitable space $V$. We are particularly interested in settings where the energy functional may depend on a parameter $\boldsymbol{p}\in D\subseteq\mathbb{R}^{d}$, typically through boundary conditions, coefficients, or right-hand sides. For a given set of parameters $\boldsymbol{p}$, the solution $u\in V$ to the PDE is given by the minimization problem
$u=\operatorname*{\arg\min}_{v\in V}E(v)$ (2.1)
where we emphasize that $E(v)$ depends on $\boldsymbol{p}$.
In this paper, we are interested in the solution operator to (2.1) depending on the parameter $\boldsymbol{p}$. Let $\mathcal{A}:D\to V$ denote this solution operator such that $u=\mathcal{A}(\boldsymbol{p})$. Here $\boldsymbol{p}$ is referred to as the problem data variable, or simply the parameter, and $D$ the parameter space. Furthermore, we assume that $\boldsymbol{p}$ is a random variable with a given probability distribution $\mathcal{P}$ over $D$.
2.1 Finite element discretization
To be able to learn the solution operator we first need to discretize $V$. For this purpose, we introduce a finite element discretization of the solution space $V$. Let $\mathcal{T}_{h}$ denote a mesh based on a finite subdivision of $\Omega$ into closed and convex elements of maximal diameter $h$, i.e., $\mathrm{diam}(T)\leq h$ for $T\in\mathcal{T}_{h}$. Let $V_{h}\subseteq V$ be suitable finite element space based on the mesh $\mathcal{T}_{h}$. The corresponding approximation $u_{h}\in V_{h}$ is attained by minimizing the energy functional over the finite element space $V_{h}$
$u_{h}=\operatorname*{\arg\min}_{v\in V_{h}}E(v)$ (2.2)
We now define the discrete version of $\mathcal{A}$,
$\mathcal{A}_{h}:D\to V_{h}\quad\text{such that}\quad u_{h}=\mathcal{A}_{h}(\boldsymbol{p})$ (2.3)
The aim is set up a neural network that approximates this mapping.
Remark 2.1.
We emphasize that the definition of $\mathcal{T}_{h}$ and $V_{h}$ are intentionally very general. The numerical examples in Section 4 will be based on first order Lagrangian (P1) finite elements on triangles. However, in the current framework, it is possible to use many other types of elements and also higher order degree polynomials.
3 Operator learning with energy minimization
3.1 Network architecture and loss function
We consider a simple feedforward and fully connected network, often referred to as multilayer perceptron. This choice of architecture is made since, to the best of our knowledge, no other work exists that combines the various components (MLP-approximated finite element solution operator, energy-based loss function, etc.) as is done here. We therefore think it is reasonable to first study and explore the basic standard architecture in this setting. Although basic, MLPs can posses good approximation capabilities as presented in the well-known work [15].
The input layer of the network is the problem data variable $\boldsymbol{p}\in D$ and the output layer is the nodal values of a finite element function in $V_{h}$. The size of the input layer is $d$, since $\boldsymbol{p}\subseteq\mathbb{R}^{d}$, and the size of the output layer equals the number of free nodes $n$ in the mesh $\mathcal{T}_{h}$, i.e., nodes that do not correspond to a Dirichlet boundary condition.
The activation function mapping the hidden layers is set to the exponential linear unit (ELU) function
$\sigma(x)=\begin{cases}x\quad&\text{if }x>0\\ e^{x}-1&\text{if }x\leq 0\end{cases}$ (3.1)
Let $\theta$ denote the trainable parameters (weights and biases) in the network and let
$\mathcal{A}_{h,\theta}:D\to V_{h}\quad\text{such that}\quad u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$ (3.2)
describe the neural network approximation of $\mathcal{A}_{h}$ in (2.3).
To train the network we use the the following energy-based loss function
$\mathcal{L}(\theta):=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right]$ (3.3)
The network thus learns the operator by minimizing the expected value of the energy with respect to the parameter $\boldsymbol{p}$. Since minimizing the energy (for a fixed $\boldsymbol{p}$) is equivalent to solving the finite element problem, we expect the network to accurately approximate the (discretized) solution operator. This is further motivated in Section 3.2. We remark that not all PDEs have a corresponding energy functional. For those that do not, one could instead use the final element residual during training, an example of weak residual minimization which is considered in [18, 5]. However, if available, the energy should be preferred to the residual. This is because the energy functional seems to be computationally cheaper to assemble and that it decomposes more easily into its local contributions, something that we exploit in Section 3.3.
3.2 Approximation error estimate
The method is based on the idea of learning finite element solution operators. This turns out to be beneficial when studying the error of the method. First we observe that during training, the loss function $\mathcal{L}(\theta)=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right]$ is indirectly minimized over a collection of functions in $V_{h}$ and $\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h}(\boldsymbol{p}))\right]$ is the minimum over the set $V_{h}$. Hence, by assuming well-trainedness of the network we may assume that the error is small, i.e.,
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))-E(\mathcal{A}_{h}(\boldsymbol{p}))\right]\leq\epsilon$ (3.4)
[…]
$E(v)=\frac{1}{2}(A\nabla v,\nabla v)-(f,v)$ (3.5)
which is the case for a Poisson problem with coefficient $A$ and right-hand side $f$. Here we consider both linear and nonlinear energies by letting $A$ depend on either the coordinates of the solution domain, like $v$, or depend on $v$ itself, respectively. We also assume the that discretization is based on Lagrangian finite elements of order $p$. The error is measured in the $H^{1}$-seminorm also informally called the $H_{0}^{1}$-norm
$|v|_{H^{1}}=\|v\|_{H_{0}^{1}}=\|\nabla v\|$ (3.6)
To obtain the approximation error estimate we will need the following results that connect the $H_{0}^{1}$-norm with the energy.
Lemma 3.1 (An identity for linear energy).
Let $E(v)=\frac{1}{2}a(v,v)-l(v)$, where $a$ is a bounded and coercive bilinear form on $V_{h}$ and $l$ is a bounded linear form on the same. Also let $a(u_{h},v)=l(v)$ for all $v\in V_{h}$, and $\|v\|_{a}^{2}:=a(v,v)$. Then
$\frac{1}{2}\|u_{h}-v\|_{a}^{2}=E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.7)
  • Proof.The left-hand side is
    $\frac{1}{2}\|u_{h}-v\|_{a}^{2}$ $=\frac{1}{2}\|v\|_{a}^{2}+\frac{1}{2}\|u_{h}\|_{a}^{2}-a(u_{h},v)$ (3.8)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+\|u_{h}\|_{a}^{2}-a(u_{h},v)$ (3.9)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+a(u_{h},\underbrace{u_{h}-v}_{\in V_{h}})$ (3.10)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+l(u_{h}-v)$ (3.11)
    $=\bigg{(}\frac{1}{2}\|v\|_{a}^{2}-l(v)\bigg{)}-\bigg{(}\frac{1}{2}\|u_{h}\|_{a}^{2}-l(u_{h})\bigg{)}$ (3.12)
    $=E(v)-E(u_{h})$ (3.13)
Under appropriate assumptions, the energy $E$ in (3.5) fulfills the requirements in Lemma 3.1 with $\|v\|_{a}^{2}=(A\nabla v,\nabla v)$. Thus applying Lemma 3.1 and noting that $\|v\|_{a}^{2}\geq\alpha\|\nabla v\|^{2}$ where $A\geq\alpha>0$, we have
$E(v)-E(u_{h})=\frac{1}{2}\|u_{h}-v\|_{a}^{2}\geq\frac{\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}$ (3.14)
We summarize this result in the following corollary.
Corollary 3.1 (An inequality for linear energy).
Let $E$ be defined by (3.5) with $A\geq\alpha>0$ and $f\in L^{2}(\Omega)$. Then there exists a positive constant such that
$\|\nabla u_{h}-\nabla v\|_{L^{2}(\Omega)}^{2}\lesssim E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.15)
Lemma 3.2 (An inequality for nonlinear energy).
Let $E(v)=\frac{1}{2}(A(v)\nabla v,\nabla v)=\frac{1}{2}\|\sqrt{A(v)}\nabla v\|^{2}$, i.e., $f=0$. Let $u_{h}\in V_{h}$ denote a minimizer to $E$, i.e., Gateaux derivative equal to zero for all $v\in V_{h}$. Assume that $A(v)\geq\alpha>0$ for all $v\in V_{h}$, and $A^{\prime}(u_{h})u_{h}\geq 0$. Then there exists a positive constant such that
$\|\nabla u_{h}-\nabla v\|_{L^{2}(\Omega)}^{2}\lesssim E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.16)
  • Proof.We start by scaling the left-hand side
    $\frac{\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}$ $=\frac{\alpha}{2}\|\nabla v\|^{2}+\frac{\alpha}{2}\|\nabla u_{h}\|^{2}-\alpha(\nabla u_{h},\nabla v)$ (3.17)
    $\leq\frac{1}{2}\|\sqrt{A(v)}\nabla v\|^{2}+\frac{1}{2}\|\sqrt{A(u_{h})}\nabla u_{h}\|^{2}-\alpha(\nabla u_{h},\nabla v)$ (3.18)
    $=E(v)+E(u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.19)
    $=E(v)-E(u_{h})+\underbrace{2E(u_{h})-\alpha(\nabla u_{h},\nabla v)}_{=R}$ (3.20)
    We consider the rest $R$ separately. For its treatment we use that one can write $A(v)=\alpha+a(v)$ for some function $a$ of $v$.
    $R$ $=2E(u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.21)
    $=(A(u_{h})\nabla u_{h},\nabla u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.22)
    $=(\alpha+a(u_{h})\nabla u_{h},\nabla u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.23)
    $=(\alpha\nabla u_{h},\nabla u_{h}-\nabla v)+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.24)
    $\leq\alpha\|\nabla u_{h}\|\|\nabla u_{h}-\nabla v\|+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.25)
    $\leq\frac{\alpha}{2\varepsilon}\|\nabla u_{h}\|^{2}+\frac{\varepsilon\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.26)
    We take $\varepsilon=1/2$ and kickback the second term. What remains is
    $\alpha\|\nabla u_{h}\|^{2}+(a(u_{h})\nabla u_{h},\nabla u_{h})=(A(u_{h})\nabla u_{h},\nabla u_{h})$ (3.27)
[…]
    =\;= 00 0 (3.29)
    Here we have used the assumption $A^{\prime}(u_{h})u_{h}\geq 0$ to get the inequality. The final step is noting that the obtained quantity is the left-hand side of the corresponding variational formulation (with test function $v=u_{h}$) which is equal to zero since $f=0$. Equivalently, the quantity is the Gateaux derivative of $E$ at $u_{h}$ (in the direction of $u_{h}$) which is equal to zero in any direction since $u_{h}$ is a minimizer to $E$. Summing up we have
    $\frac{\alpha}{4}\|\nabla u_{h}-\nabla v\|^{2}\leq E(v)-E(u_{h})$ (3.30)
    which is the desired inequality.
Under these assumptions on the energy and the discretization we derive the following error bound.
Theorem 3.1 (An estimate for the approximation error).
For parameters $\boldsymbol{p}$ from a probability distribution $\mathcal{P}$, let $\mathcal{A}$ denote the solution operator to the continuous problem (2.1). Given a finite element discretization of this problem with mesh size $h$ and based on Lagrangian elements of polynomial degree $p$, let $\mathcal{A}_{h,\theta}$ denote the corresponding neural network approximation of the discrete solution operator $\mathcal{A}_{h}$. If we assume well-trainedness of the network with error bound $\epsilon>0$, i.e., (3.4), then there exists a positive constant such that
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|\nabla\mathcal{A}(\boldsymbol{p})-\nabla\mathcal{A}_{h,\theta}(\boldsymbol{p})\|_{L^{2}(\Omega)}^{2}\right]\lesssim h^{2p}\bigg{(}\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|D^{p+1}\mathcal{A}(\boldsymbol{p})\|_{L^{2}(\Omega)}^{2}\right]\bigg{)}+\epsilon$ (3.31)
  • Proof.Recall that $u=\mathcal{A}(\boldsymbol{p})$ and $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$. Hence, the argument to the expectation operator on left-hand side is
    $\|\nabla u-\nabla u_{h,\theta}\|^{2}$ $\lesssim\|\nabla u-\nabla u_{h}\|^{2}+\|\nabla u_{h}-\nabla u_{h,\theta}\|^{2}$ (3.32)
    $\lesssim h^{2p}\|D^{p+1}u\|^{2}+E(u_{h,\theta})-E(u_{h})$ (3.33)
    Here we have started with an error split using ±$\pm u_{h}$. Then we have applied a standard error estimate for the finite element error and a result like Corollary 3.1 or Lemma 3.2 for the learning error where we have used that $u_{h,\theta}\in V_{h}$. Taking the expected value of both sides preserves the inequality. Using linearity of the expected value and the well-trainedness assumption (3.4) of the network gives the desired result. ∎
3.3 Efficient learning
The network will be trained using an iterative method based on the stochastic gradient descent (SGD), e.g., the Adam optimizer. To train the network efficiently, the computation of the gradient of the loss function in (3.3) needs to be fast. Typically, a mini-batch of instances is used to evaluate the expected value in (3.3). In each iteration of the optimization algorithm, a sample of size $M$, i.e., $\{\boldsymbol{p}_{i}\}_{i=1}^{M}$, is generated from the distribution given by $\mathcal{P}$. The loss function in (3.3) is thus approximated by
$\frac{1}{M}\sum_{i=1}^{M}E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.34)
[…]
$E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))=\sum_{T\in\mathcal{T}_{h}}E_{T}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.35)
where $E_{T}$ denotes the local energy contribution on the element $T$. For, e.g,. a Poisson problem, $E_{T}$ would require the local assembly of the stiffness and load terms. See Section 4 for further examples in practice. We note that these computations can be done in parallel on, e.g., a GPU. For large problems with many elements, we suggest to also use mini-batches of elements. The mini-batch is based on a uniform selection of $N$ elements $\{T_{i}\}_{i=1}^{N}$ in each iteration, which is similar to the use of mini-batches in SGD. This means that the loss function is approximated by
$\frac{1}{M}\sum_{i=1}^{M}\sum_{j=1}^{N}E_{T_{j}}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.36)
Note that we do not divide by $N$, since we do not aim to approximate an expected value but rather the integral over the full domain $\Omega$.
4 Examples
In this section we provide three examples to evaluate how the proposed method performs in practice. The first example considers a relatively simple parameterized PDE to introduce the energy minimization approach. The second example concerns a random PDE with a coefficient depending on a Gaussian random field. Gaussian random fields are of particular interest in, e.g., geophysical applications where the field can be used to describe porous media. The trained network can be used in uncertainty quantification, where many different solutions to the given PDE, corresponding to different realizations of the field, are needed. The third example concerns an application of learning solution operators for nonlinear elasticity problems. Here, the network output is combined with conventional finite element software to preserve accuracy of the final result and at the same time speed up the computations. This is done by using the learned solution as an initial guess for Newton’s method.
The implementation used for the examples is based on the code presented in [34] which is publicly available at https://github.com/nmwsharp/neural-physics-subspaces. The training of the neural networks in the examples was enabled by resources provided by National Academic Infrastructure for Supercomputing in Sweden (NAISS) [1]. More precisely, the training has been performed on NVIDIA Tesla A100 HGX GPUs on the Alvis cluster.
4.1 A parameterized PDE
Consider the following parameterized Poisson problem
$-\nabla(A(\boldsymbol{p})\cdot\nabla u)$ $=f\quad$ in $\text{in }\Omega$ (4.1)
$u$ $=0$ on $\text{on }\partial\Omega$ (4.2)
with $\boldsymbol{p}=(x_{0},y_{0},r)$, $f=1$, and
$A(\boldsymbol{p},x,y)=A(x_{0},y_{0},r,x,y)=0.1+\exp(-((x-x_{0})^{2}+(y-y_{0})^{2})/r)$ (4.3)
This is a coefficient which is equal to $0.1$ everywhere, except for a spike given by the exponential term. The parameters $\boldsymbol{p}=(x_{0},y_{0},r)$ determine the location and the radius of the spike. In this example we consider $\Omega=[-1,1]\times[-1,1]$, a uniform distribution for the coordinates $x_{0},y_{0}\sim\mathcal{U}([-1,1])$ (independent), and another uniform distribution for the radius $r\sim\mathcal{U}([0.01,0.2])$.
The weak form of (4.1)-(4.2) is given by the following: Find $u\in H^{1}_{0}$ such that
$(A(\boldsymbol{p})\nabla u,\nabla v)_{L^{2}}=(f,v)_{L^{2}}\quad\forall v\in H^{1}_{0}$ (4.4)
which induces the following energy
$E(v)=\frac{1}{2}(A(\boldsymbol{p})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}}$ (4.5)
For the discretization of $H^{1}_{0}$ we consider classical piecewise linear and continuous finite elements based on a triangulation $\mathcal{T}_{h}$ of the domain $\Omega$. Let $V_{h}$ denote this finite element space. Note that for $v\in V_{h}$ the computation of the energy requires the stiffness and load terms, which can be assembled locally on each triangle in the domain, i.e.,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{p})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}$ (4.6)
We consider a uniform mesh of size $h=\sqrt{2}\cdot 2^{-4}$, which results in 961 degrees of freedom. This means that the output layer of the network contains 961 nodes, while the input layer consists of 3 nodes corresponding to $(x_{0},y_{0},r)$. The network is trained using the Adam optimizer and $10^{6}$ iterations. We use two different widths to explore the impact on accuracy. The training times are reported in Table 1. We deduce that using batches of triangles ($T=32$) has no effect. The training times are roughly the same as for the full mesh. This is due to the fact that energy computation is done in parallel and the problem is relatively small (the GPU is not maximized).
Table 1: Training and inference time for the parameterized problem in Section 4.1. Output layer of size 961. Trained on an A100 GPU.
$T$ Dimension Training time Inference time
$32$ $4\cdot 256$ 458 s 1 ms
Full $4\cdot 256$ 433 s 1 ms
$32$ $4\cdot 512$ 440 s 1 ms
Full $4\cdot 512$ 426 s 1 ms
To investigate the error of the method we compute (an approximation of) the expected value of the difference between the output of the network and the finite element solution in different norms. Given a sample of parameters $\{\boldsymbol{p}_{i}\}_{i=1}^{K}$ we compute the difference in the energy, the $L^{2}$-norm, and the $H^{1}$-norm, i.e., we compute the relative errors
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|$ (4.7)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})-\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{L^{2}}/\|\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{L^{2}}$ (4.8)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})-\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{H^{1}}/\|\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{H^{1}}$ (4.9)
Note that the difference in the energy (4.7) gives an approximation of the relative well-trainedness of the network, cf., (3.4). In Table 2 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{p}$. Corresponding histogram plots are shown in Figure 2 for the network of size $4\cdot 256$ and batches using all elements in the mesh. We note that the error distribution is small which implies that the network is well-trained over the distribution of the parameters. Furthermore, comparing the errors in Table 2 we see that increasing the width of the network improves the approximation properties. We also note that the error naturally increases when using smaller batch sizes of triangles.
Table 2: Relative errors for the parameterized problem in Section 4.1. Output layer of size 961. Sample size $K=10^{4}$.
$T$ Dimension Energy Std $L^{2}$ Std $H^{1}$ Std
$32$ $4\cdot 256$ 0.0061 0.0055 0.026 0.011 0.065 0.026
Full $4\cdot 256$ 4.1e-4 0.0013 0.0036 0.0041 0.014 0.013
$32$ $4\cdot 512$ 0.0045 0.0042 0.024 0.010 0.057 0.021
Full $4\cdot 512$ 2.0e-4 7.3e-4 0.0020 0.0022 0.0092 0.0093
(a) Energy
(b) $L^{2}$-norm
(c) $H^{1}$-norm
Figure 2: Histogram plots for the different relative errors for a network of size $4\cdot 256$ using all elements (full) per batch, i.e., the network in the second row of Table 2. Sample size $K=10^{4}$.
To show a case where batches of triangles does have an effect on the training times, we consider a finer mesh of size $h=\sqrt{2}\cdot 2^{-6}$ which results in 16129 degrees of freedom and 32768 triangles. We also increase the width of the neural network to $4\cdot 2048$. In Table 34 the results are reported for batches of size 3277 ($10\%$ of the elements) compared to the full mesh. We observe that the training time is reduced by using a smaller batch of triangles, since the GPU is now maximized and the energy computation can not be completely parallelized in each iteration. The learning error increases though, since the number of iterations is the same in both cases. We emphasize that for large problems and the choice of GPU model, the batches can have significant impact on the computational times. On a laptop GPU, this could also be crucial to be able to train the network at all.
Table 3: A neural network of $4\cdot 2048$ hidden layers for the parameterized problem in Section 4.1. Output layer of size 16129. Trained on an A100 GPU.
$T$ Training time Inference time
$3277$ 1720 s 1 ms
Full 1962 s 1 ms
Table 4: Relative errors for the parameterized problem in Section 4.1. Output layer of size 16129. Sample size $K=10^{4}$
$T$ Dimension Energy Std $L^{2}$ Std $H^{1}$ Std
$3277$ $4\cdot 2048$ 0.0033 0.0014 0.0090 0.0042 0.050 0.01
Full $4\cdot 2048$ 1.2e-04 2.7e-04 0.0026 0.0013 0.0082 0.0050
We end this numerical example by noting two things: Firstly we note that the inference time, i.e., the time for one forward pass through the network is very small, around 1 ms in all cases studied. This shows the strength of learning the solution operator. Once trained, the solution for a given parameter is obtained using very little effort. Secondly we note that the relative energy learning error (4.7) (“Energy” in Table 2 and 4) decreases with increased network width, both for hidden layers and output layer (an increase here means a decrease in the mesh size $h$). This lends credence to the well-trainedness assumption (3.4), and thus to Theorem 3.1, where it is used.
4.2 Gaussian random fields
We consider a Poisson problem with a random coefficient based on a Gaussian random field
$-\nabla(A(\omega)\cdot\nabla u)$ $=f\quad$ in $\text{in }\Omega$ (4.10)
$u$ $=0$ on $\text{on }\partial\Omega$ (4.11)
where $A(\omega)$ is defined as $\exp{a(\omega)}$ where $a(\omega)$ is a Gaussian random field with covariance function given by
$C(x,y)=\exp{\bigg{(}-\frac{|x-y|^{2}}{2L^{2}}\bigg{)}}$ (4.12)
where $L$ is the correlation length scale. We use a Karhunen-Loève expansion to approximate the field
$a(x,y,\omega)\approx\mu+\sigma\sum_{i=1}^{N_{KL}}\sqrt{\lambda_{i}}\phi_{i}\xi_{i}(\omega)$ (4.13)
where $\mu$ and $\sigma$ are the constant mean and standard deviation of the field, respectively, $\lambda_{i}$ and $\phi_{i}$ are the eigenvalues and eigenvectors corresponding to the covariance, respectively, and $\xi_{i}\sim\mathcal{N}(0,1)$ are i.i.d random variables. The number $N_{KL}$ determines the number of terms in the truncated sum. Similarly to the Poisson problem in Section 4.1, the energy in this setting is given by
$E(v)=\frac{1}{2}(A(\boldsymbol{\xi})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}}$ (4.14)
where $\boldsymbol{\xi}(\omega)=(\xi_{1}(\omega),...,\xi_{N_{KL}}(\omega))$. The vector $\boldsymbol{\xi}$ takes different values depending on the realization $\omega$ and is thus considered the parameter in this case. For the local computations on a triangle we have, c.f., Section 4.1,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{\xi})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}$ (4.15)
For the numerical experiment we choose $\mu=0$, $\sigma=1$, $L=0.5$, and $N_{KL}=9$. For the discretization, we consider a uniform mesh of size $h=\sqrt{2}\cdot 2^{-5}$, which results in 3969 degrees of freedom. This means that the output layer of the network contains 3969 nodes, while the input layer consists of 9 nodes corresponding to the length of the vector $\boldsymbol{\xi}$. We emphasize that given a realization $\omega$, the trained neural network approximates the solution operator by $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{\xi}(\omega))$. In Table 5 we report training time for a $4\cdot 512$ network. We do not consider batches of triangles in this example, since the number of degrees of freedom is relatively small.
Table 5: A neural network of $4\cdot 512$ hidden layers for the Gaussian random field problem in Section 4.2. The columns show the total training time and inference time. Trained on an A100 GPU.
T Training time Inference time
Full 426 s 1 ms
Similar to first example in Section 4.1 we investigate the error by computing an approximation of the expected value of the difference between the output of the network and the finite element solution. That is, given a sample of parameters $\{\boldsymbol{\xi}_{i}\}_{i=1}^{K}$ we compute the relative errors
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|$ (4.16)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i})-\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{L^{2}}/\|\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{L^{2}}$ (4.17)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i})-\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{H^{1}}/\|\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{H^{1}}$ (4.18)
In Table 6 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{\xi}$. Corresponding histogram plots are shown in Figure 3. We note that error distribution is small which implies that the network is well-trained over the distribution of the parameters.
Table 6: A neural network of $4\cdot 512$ hidden layers for the Gaussian random field problem. The columns show relative errors in different norms using all elements in the mesh. Sample size $K=10^{4}$.
Energy $E(\cdot)$ $L^{2}$-norm $H^{1}$-norm
Mean 3.2e-4 0.0052 0.0131
Std 4.8e-4 0.0029 0.0064
(a) Energy
(b) $L^{2}$-norm
(c) $H^{1}$-norm
Figure 3: Histogram plots for the different relative errors for a network of size $4\cdot 512$ using all elements (full) per batch, i.e., the network in Table 6. Sample size $K=10^{4}$.
For random PDEs we are typically interested in a quantity of interest (QoI), for instance the expected value of the $L^{2}$-norm of the solution
$\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right]$ (4.19)
or the average point value at some coordinate $x_{0}$
$\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right]$ (4.20)
A classical Monte Carlo approach to approximate the quantity will require many evaluations of the solution operator $u=\mathcal{A}(\boldsymbol{\xi})$. For the finite element approximation, this means that we would need to compute an approximation for each realization $\omega$. In particular, the stiffness matrix needs to be reassembled for each such realization. However, for the trained neural network, retrieving an approximation for a given realization is a simple forward pass of the network which is typically very fast (around 1 ms). This step is also easily parallelized on the GPU using the JAX framework. To show the advantage of the neural network approach, we compare the computational times for evaluating the quantities of interest (4.19) and (4.20) using the trained neural network with the corresponding times from using a finite element approach with the same degrees of freedom. The implementation is done using the FEniCSx library [4]. The GPU computations are performed on the same A100 GPU used for the training and the CPU computations on an Intel(R) Core(TM) i9-11950H 2.60GHz. The results are presented in Table 78.
Table 7: Times for computing $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right]$ using $10^{4}$ samples.
Method QoI GPU (parallel) CPU (sequential)
NN 0.186 0.21 s 25 s
FEM 0.181 79 s
Table 8: Times for computing $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right]$ for $x_{0}=(0,0)$ using $10^{4}$ samples.
Method QoI GPU (parallel) CPU (sequential)
NN 0.311 0.26 s 20 s
FEM 0.310 71 s
We observe that the neural network clearly outperforms the FEM implementation. In particular when the inference is parallelized on the GPU, but also on a classical CPU when the inferences are done sequentially. This is thus an example of a setting which really shows the potential strength of learning the solution operator using neural networks.
4.3 Nonlinear elasticity
A promising application of solution operator networks is to use the output as an initial guess for Newton’s method when solving nonlinear problems, see, e.g., [16, 28, 2]. This is a complementary way of using machine learning for PDE solving rather than as a complete solver method, i.e., to consider the operator network’s prediction the finished product. Fair criticism towards the latter includes potential loss of accuracy when compared to standard methods such as FEM. Using the network prediction as initial guess for Newton’s method is a way to preserve accuracy while at the same time potentially speed up the iterative process. It is also a natural example of when machine learning frameworks can be combined with existing software for standard methods, thus motivating the learning of the standard method’s approximation, here the final element solution. The finite element software we use here is the popular finite element library FEniCS [23].
Using neural networks to predict hyperelasticity solutions has been done in, e.g., [27] with a discretization-free method and in [7] with a meshfree collocation method. Here we consider a nonlinear neo-Hookean elasticity problem: a two dimensional cantilever beam under the influence of a given external force. The external force is defined by three variables: the midpoint $p$ of a fixed-length surface interval of application on top of the beam and the horizontal and vertical magnitudes $F_{x}$ and $F_{y}$, respectively. The solution to the problem is the displacement field over the beam. We generate a structured uniform triangular 40x2 mesh of the resting beam and consider standard P1 finite elements. This results in 240 DOFs, since every mesh node ($41\cdot 3=123$) has two DOFs (one for each dimension) and the DOFs of the left-most three nodes are removed since that is where the beam is fixed.
The solution operator we aim to learn is thus the map from the 3 force variables to the 240 finite element DOFs giving the corresponding displacement of the beam. We use a simple MLP architecture that takes 3 input variables (the force variables), has 4 hidden layers of width 256 with ELU activation and one output layer of width 240 (number of DOFs) with no activation. Just as in the previous examples we use the energy functional corresponding to the PDE as the foundation for the loss function. Here, the energy is
$E(v)=\int_{\Omega}\Psi(v)\,\text{dx}-\int_{\partial\Omega\cap\text{supp}(F)_{p}}F\cdot v\,\text{ds}$ (4.21)
where $\Psi$ is the neo-Hookean elastic stored energy density and $F=(F_{x},F_{y})$ is the external force field. Gravity is not included.
During training, we pick 32 randomly selected tuples of force variables in each iteration to compute an average loss. The midpoint of the interval of application is sampled uniformly along the beam $p\sim\mathcal{U}=\mathcal{U}([x_{\min},x_{\max}])$, where $x_{\min}$ and $x_{\max}$ are the left and right endpoints of the beam, respectively. The horizontal and vertical magnitudes are sampled from normal distributions with mean 0 and standard deviation $\sigma$, i.e., $F_{x},F_{y}\sim\mathcal{N}(0,\sigma^{2})$. The optimization is performed with the Adam optimizer where we perform $10^{6}$ iterations with a decreasing learning rate. The learning rate starts at 1e-4 and after every 250k iterations it is decreased by a factor of 0.5.
Sampling the force magnitudes from normal distributions with mean 0 means that we train the networks more for problems with solutions close to the zero function, i.e., the unbent beam. It is thus reasonable to expect the networks to perform better for such problems than for problems with large forces resulting in large deformations. The previously used relative learning errors are thus unsuitable here, since they could potentially give very large errors for problems with solutions close to zero and smaller errors for more deformed solutions. Letting $\boldsymbol{F}=(p,F_{x},F_{y})$ and $|\Omega|$ denote the Lebesgue measure of $\Omega$, we therefore instead compute the $\Omega$-relative learning errors
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{F}_{i}))|/|\Omega|$ (4.22)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i})-\mathcal{A}_{h}(\boldsymbol{F}_{i})\|_{L^{2}}/|\Omega|^{1/2}$ (4.23)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i})-\mathcal{A}_{h}(\boldsymbol{F}_{i})\|_{H_{0}^{1}}/|\Omega|^{1/2}$ (4.24)
We consider two different force situations: basic beam bending and extreme bending.
4.3.1 Basic bending
Here, we only consider one simple sampling case for the force variables in which we do not include a horizontal magnitude. The force variables are sampled by
$p\sim\mathcal{U}\quad F_{x}=0\quad F_{y}\sim\mathcal{N}(0,1)$ (4.25)
Training and inference times for the basic bending case on both an A100 GPU and an Apple M1 CPU are presented in Table 9. The training times clearly show the benefit of GPU training.
Table 9: Training and inference times for the basic bending case.
Hardware Training time Inference time
GPU (A100) 558 s 0.8 ms
CPU (M1) 2393 s 0.13 ms
Learning errors for the basic bending case are presented in Table 10 and corresponding histogram plots are presented in Figure 4. The histograms show that the error distributions are skewed towards zero, indicating some degree of well-trainedness.
Table 10: $\Omega$-relative learning errors in different norms for the basic bending case. Sample size $K=10^{3}$.
Training & test data Measure Energy $E(\cdot)$ $L^{2}$-norm $H_{0}^{1}$-norm
$p\sim\mathcal{U}\quad F_{x}=0\quad F_{y}\sim\mathcal{N}(0,1)$ Mean 0.0058 0.0115 0.0477
Std 0.0085 0.0126 0.0481
(a) Energy
(b) $L^{2}$-norm
(c) $H_{0}^{1}$-norm
Figure 4: Histogram plots for the different $\Omega$-relative learning errors for the basic bending case, i.e., Table 10. Sample size $K=10^{3}$.
We now consider the specific problem of computing the displacement when the external force is located at the right end of the beam and directed downwards with magnitude 1.0. We use the operator network for the basic bending case and perform the Newton iterations with a FEniCS implementation of the problem. All computations are performed on an Apple M1 CPU. Results from using the network prediction and a standard initial guess (the zero function) as input for Newton’s method are presented in Table 11 and Figure 5.
Table 11: Computational times on an Apple M1 CPU for using different initial guesses for Newton’s method.
Initial guess $E(u_{h})$ Inference time Newton solver time Total time
Network prediction -0.014397 9.4e-5 s 0.015678 s 0.015772 s
Zero function -0.014397 0.022658 s 0.022658 s
(a) Network prediction as initial guess.
(b) Zero function as initial guess.
Figure 5: Different initial guesses for Newton’s method with resulting converged solutions for a basic bending case. The converged solution $u_{h}$ is the same in both cases.
The results in Table 11 and Figure 5 show that using the network prediction as initial guess for Newton’s method preserves accuracy while at the same time provides a speed-up compared to using the standard zero function as initial guess. However, to properly benefit from using the network prediction as initial guess, a large enough number of problems will have to be solved so that all individual problem gains compensate for the training time. From Table 9, we compute that one needs to solve at least 81,034 problems with GPU training and 347,517 with CPU training to get this benefit. This might sound like a lot, but if the trained network is incorporated into software intended for many users or if many instances have to be solved, e.g., in an optimization process or for a time-dependent version of the problem, the numbers quickly diminish in perspective.
4.3.2 Extreme bending
Here, we consider three more advanced sampling cases for the force variables:
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ (4.26)
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ (4.27)
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ (4.28)
Training and inference times for the extreme bending cases on both an A100 GPU and an Apple M1 CPU are presented in Table 12. Again the results show the benefits of GPU training.
Table 12: Training and inference times for the extreme bending cases. The A100 GPU Utilization was 46% in all three cases.
Training & test data Hardware Training time Inference time
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ GPU (A100) 583 s 0.85 ms
CPU (M1) 2398 s 0.17 ms
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ GPU (A100) 583 s 0.88 ms
CPU (M1) 2438 s 0.14 ms
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ GPU (A100) 582 s 0.86 ms
CPU (M1) 2672 s 0.13 ms
Learning errors for the extreme bending cases are presented in Table 13 and histograms for the last case are presented in Figure 6. The learning errors increase with increased variance of the data distributions. This is reasonable since wider distributions can be thought of as larger domains of the map that the network approximates. Larger domains, mean more complicated maps and in turn harder to train networks that are harder to train well. Nevertheless, the histograms for the last case all show small error distributions skewed towards zero, indicating some degree of well-trainedness.
Table 13: $\Omega$-relative learning errors in different norms for the extreme bending cases. Sample size $K=10^{3}$.
Training & test data Measure Energy $E(\cdot)$ $L^{2}$-norm $H_{0}^{1}$-norm
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ Mean 0.0443 0.0395 0.1564
Std 0.1720 0.0749 0.2928
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ Mean 0.1396 0.0574 0.2236
Std 0.4489 0.1000 0.3819
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ Mean 0.2299 0.0704 0.2669
Std 0.7610 0.1104 0.4052
(a) Energy
(b) $L^{2}$-norm
(c) $H_{0}^{1}$-norm
Figure 6: Histogram plots for the different $\Omega$-relative learning errors for the last extreme bending case, i.e., the last row of Table 13. Sample size $K=10^{3}$.
We now consider the specific problem of computing the displacement when the external force is again located at the right end of the beam but successively turns around the fixed end, forcing the beam to curl. This is different from the basic bending example since the force now changes when the beam has been bent to a certain degree. Doing this with the standard approach means performing sets of Newton iterations: From an initial state of the beam and an initial force, Newton iterations are performed until convergence. The converged state is then used as the new starting state together with a new force, and so on. The last converged state is the desired one and the other ones are simply intermediates. The forces we consider for the standard approach are:
$F_{1}=(0.0,-0.99)\quad F_{2}=(-1.5,0.0)\quad F_{3}=(0.0,5.0)$ (4.29)
This means that we will get three states: two intermediates and one final curled state. For the machine learning approach, we note that the operator networks we have trained cannot reach the final curled state of the beam. This is because the forces used as input to the network are all applied to the unbent beam. However, the operator networks can still provide predictions of intermediate states that can be useful as initial guesses. We use an operator network from the last extreme bending case, i.e., where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training, and take the network intermediate prediction coming from $F=(-10.0,-5.0)$ as initial guess. We again perform the Newton iterations with a FEniCS implementation of the problem. All computations are performed on an Apple M1 CPU. Results from using the network intermediate state and the successive beam configurations starting from the standard initial guess (the zero function) as input for Newton’s method are presented in Table 14 and Figure 7.
Table 14: Computational times on an Apple M1 CPU for using different initial guesses for Newton’s method. The network intermediate state comes from an operator network where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training.
Initial guess $E(u_{h})$ Inference time Newton solver time Total time
Network intermediate 0.067984 9.5e-5 s 0.019925 s 0.02002 s
Zero function 0.067984 0.04784 s 0.04784 s
(a) Network intermediate as initial guess. The unbent beam is present for reference.
(b) Zero function (unbent beam) as initial guess. The two intermediate states are also present.
Figure 7: Different initial guesses for Newton’s method with resulting converged solutions for an extreme bending case. The converged solution $u_{h}$ is the same in both cases. The network intermediate state comes from an operator network where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training.
The results in Table 14 and Figure 7 show that just as in the basic bending case, using the network output as initial guess for Newton’s method preserves accuracy while at the same time provides a speed-up compared to using the standard zero function as initial guess. Furthermore, the gain from using the network output is higher in the extreme bending case compared to the basic one. This is easily seen by comparing the total times in Table 11 and 14. For the basic bending case we have that the total time with the network output is about 70% of the total time with the zero function. The corresponding quantity for the extreme bending case is 42%. This higher gain naturally affects the number of problems that have to be solved in order to compensate for the training time of the network. From the last case in Table 12, we compute that one needs to solve at least 20,921 problems with GPU training and 96,047 with CPU training to get a benefit from using the network output as initial guess when taking the training time into account. These numbers for the extreme bending case are 26% and 28% of those of the basic bending case for GPU and CPU training, respectively. We point out that the higher gain from using a network in an extreme bending case could in fact be even higher. This is so since here we only use the network to get an intermediate state, not a prediction of the solution as in the basic bending case. From Table 11 and 14, the Newton solver time for using a prediction in the basic bending case is 79% of the one for using an intermediate state in the extreme case, suggesting the even higher computational gain. As a final remark we point out that the two Newton solver times do not seem to differ that much while the corresponding beam configuration pairs (starting and converged states) differ quite a lot visually: From Figure 5 and 7, the starting state is much closer to the converged one in the basic bending case compared to in the extreme one. We take this to mean that most of the Newton computational work is performed towards the end, fine-tuning the beam configuration to the final state.
To summarize, combining operator networks with Newton’s method preserves accuracy and can provide a speed-up. In the case of nonlinear elasticity, the greater the deformation, the greater the potential speed-up from using operator networks.
5 Conclusions and outlook
We have presented a machine learning framework for learning solutions to a class of PDE problems. A core idea of the framework is to learn the corresponding discrete solution of some standard numerical method instead of aiming for the exact solution. The reason being that the standard method could be used to aid and enhance the framework. This core idea can in general be applied to various machine learning frameworks and standard methods but here we have considered a simple MLP-architecture together with energy minimization for the framework and FEM as the standard numerical method. We have presented both theoretical results (approximation error estimate) and practical applications (Newton’s method) that demonstrate how the framework may be beneficially combined with FEM. We have also presented pure framework results that show strengths and limitations of it as well as potential applications. These results are the learning errors, the usage of batches of elements for the energy during training, and the computation of quantities of interest.
Concerning avenues for future work, besides looking into more advanced network architectures and training algorithms, the last elasticity example provides a natural starting point. We note that in this example (the extreme bending case), the neural network is limited by the fact that the external forces are always applied to the initial state (the unbent beam). This could be improved by introducing time dependency where the forces are allowed to change during the bending process. This would also most likely mean that the network needs to take the current position of the beam as an input parameter. This is an interesting path for future research.
Acknowledgement.
This research was supported in part by the Swedish Research Council Grant No. 2021-04925 and Grant No. 2022-03543, and the Swedish Research Programme Essence.